OpenStudy (anonymous):

Solve this equation in terms of a B = 3a – 4

OpenStudy (anonymous):

@sirizarry78 Thank you for helping

OpenStudy (anonymous):

You want to isolate a. First add 4 to both sides of the equation, you get B + 4 = 3a Now divide both sides by 3, you get a = (B+4)/3

OpenStudy (anonymous):

Okay then what

OpenStudy (anonymous):

Didnt you want to solve for a? If thats what you wanted to do, then we accomplished what we wanted to do. a = (B+4)/3

OpenStudy (anonymous):

Ohh right, sorry.

OpenStudy (anonymous):

welcome.